OpenStudy (anonymous):

The perimeter of a rectangle is 100 feet. The length can be represented by (6x - 2) feet and the width can be represented by (x + 3) feet. What is the value of x? (1 points) 14.1 7 10 6.4

OpenStudy (7thgrader):

Make sure you didn't give any medals on a post yet, and open the same post in a new tab, then give a medal to two different people.

OpenStudy (anonymous):

2*(6x-2)+2*(x+3)=100 (6x-2)+(x+3)=50 7x+1=50 7x=49 x=?

OpenStudy (anonymous):

P = 2l + 2w

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Correct :)